OpenStudy (anonymous):

3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5

OpenStudy (anonymous):

Plug it in. 3(3)2-(-2)3-(-2)-(-5) Which will give you 31.

OpenStudy (anonymous):

3(3)^2 - (-2)^3 - (-2)^3 - (-5) = 27 + 8 + 8 +5 = 48